OpenStudy (anonymous):

Divide and simplify to lowest terms \[\frac{ x^{2}-64\\ }{ 5x+15 } \times \frac{ x+3 }{ x-8 }\]

OpenStudy (anonymous):

\[\frac{ a }{ b } \times \frac{ c }{ d } = \frac{ ad+bc }{ bd }\] Try that and let me know what you get :)

OpenStudy (anonymous):

\[\frac{ x+8 }{ 5 }\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Correct! Great work :)

OpenStudy (anonymous):

@sarahusher sorry you're mistaken , what you've written is for \[\frac{ a }{ b }+\frac{ c }{ d }\]

OpenStudy (anonymous):

It is! Oh dear, I apologise..sorry it should be \[\frac{ a }{ b }\times \frac{ c }{ d }=\frac{ ac }{ bd }\] Gosh I'm so sorry, @learner123

OpenStudy (anonymous):

thanks @itsmylife !

OpenStudy (anonymous):

no it seems fair, lol you might be sleepy whilst writing this, np lady :)

OpenStudy (anonymous):

I didn't even use that formula to work it out, lol no idea where my head was. They got the right answer regardless.... I can't even blame it on tiredness, it's the afternoon! lol

OpenStudy (anonymous):

whoa pretty weird ..still i wonder how come you got right answer